OpenStudy (anonymous):

Integrate: dy/(1-y^2) Anybody knows how to integrate this without the weird Arctanh(x) ?

OpenStudy (anonymous):

\[\int\limits \frac{dy}{ (1-y^2)}\] Let \[y= sinx \rightarrow dy = cosx dx\] Thus \[\int\limits \frac{dy}{ (1-y^2)} = \int\limits \frac{cosx dx}{ (1-\sin^2x)}= \int\limits \frac{cosx dx}{ (\cos^2x)}\] \[= \int\limits \frac{ dx}{ (cosx)} = \int\limits \sec x dx = \log|\sec x + \tan x| +c\]